OpenStudy (anonymous):

5. Evaluate u + xy, if u = 18, x = 10, and y = 8. (1 point)

OpenStudy (solomonzelman):

u + xy YOU KNOW THAT... u = 18, x = 10, and y = 8 PLUG THE VALUES IN.... 18+(10)(8) 18+(10)(8)= 18+80= 98

OpenStudy (anonymous):

thank you @SolomonZelman

OpenStudy (solomonzelman):

You welcome :)